What is the freezing point depression for a 2 m solution of aluminum chloride? show all steps
Answer #1

AlCl3 dissociates as
AlCl3↔Al3+ + 3Cl-

Assuming complete dissociation, i=4

Cryoscopic constant for water (Kf)=1.86 °C/m

Therefore, depression in freezing point (∆Tf)= i* Kf * m=4×1.86 °C/m ×2=14.88°C